x+2x+5=x+19 please help my daughter understand how to do this type of math she is 10 yrs old thank you
Answer by MILIAJEY(4)

X+2X+ 5= X+19
BRING ALL THE X TERMS TO ONE SIDE: AND ALL THE INTEGERS TO THE OTHER SIDE:
X+ 2X -X = 19 -5
2X = 14
X = 7